Can someone please show all your work and s

Can someone please show all your work and steps for this problem? Please write really legibly and make it easier for me to understand... The answer is ALREADY there so PLEASE make sure to double check your work and make sure your answer matches. Thanks in advance!
Answer #1

initial velocity u=7.92 m/s

the acceleration a=-[1.61i+4.72j]

the magnitude of acceleration a =4.99 m/s^2

this concept belongs motion in two dimensional

according to the equation of kinamatices

the distance S=u^2/2a

=7.92^2/2*4.99=6.3 m